OpenStudy (aaronandyson):

If sec A = sqrt2 find 3 cot^2(A) + 2 sin^2(A)/tan^2(A) - cos^2(A)

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (aaronandyson):

@Compassionate

OpenStudy (aaronandyson):

@Preetha @Palmo4ka @Somy @Hero @IrishBoy123

OpenStudy (michele_laino):

If we apply the definition of sec(A), then we can write: \[\frac{1}{{\cos \left( A \right)}} = \sqrt 2 \]

OpenStudy (aaronandyson):

cos(A) = 1/sqrt2

OpenStudy (michele_laino):

ok!

OpenStudy (aaronandyson):

Now?

OpenStudy (michele_laino):

please you can compute (sin(A))^2, using the fundamental identity, namely: \[{\left( {\sin \left( A \right)} \right)^2} = 1 - {\left( {\cos \left( A \right)} \right)^2} = ...?\]

OpenStudy (aaronandyson):

1 - 1/2?

OpenStudy (michele_laino):

yes!

OpenStudy (aaronandyson):

1/2?

OpenStudy (michele_laino):

yes!

OpenStudy (aaronandyson):

now?

OpenStudy (michele_laino):

now please use these identities, in order to find: (Tan(A))^2, and (cot(A))^2: \[\begin{gathered} {\left( {\tan \left( A \right)} \right)^2} = \frac{1}{{{{\left( {\cos \left( A \right)} \right)}^2}}} - 1 = 2 - 1 = ...? \hfill \\ {\left( {\cot \left( A \right)} \right)^2} = \frac{1}{{{{\left( {\sin \left( A \right)} \right)}^2}}} - 1 = 2 - 1 = ...? \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

1 and 1?

OpenStudy (michele_laino):

perfect!

OpenStudy (michele_laino):

finally, substitute those values, namely: \[\begin{gathered} {\left( {\sin \left( A \right)} \right)^2} = {\left( {\cos \left( A \right)} \right)^2} = \frac{1}{2} \hfill \\ {\left( {\tan \left( A \right)} \right)^2} = {\left( {\cot \left( A \right)} \right)^2} = 1 \hfill \\ \end{gathered} \] into your original expression: \[\frac{{3{{\left( {\cot \left( A \right)} \right)}^2} + 2{{\left( {\sin \left( A \right)} \right)}^2}}}{{{{\left( {\tan \left( A \right)} \right)}^2} - {{\left( {\cos \left( A \right)} \right)}^2}}} = ...?\]

OpenStudy (aaronandyson):

sin was 1/2 right?

OpenStudy (aaronandyson):

I mean sin(A) = 1/2?

OpenStudy (michele_laino):

no, please: \[\sin \left( A \right) = \frac{1}{{\sqrt 2 }}\]

OpenStudy (aaronandyson):

sorry forgot the ^2 sign and please continue

OpenStudy (michele_laino):

ok! Please what is: \[\frac{{3{{\left( {\cot \left( A \right)} \right)}^2} + 2{{\left( {\sin \left( A \right)} \right)}^2}}}{{{{\left( {\tan \left( A \right)} \right)}^2} - {{\left( {\cos \left( A \right)} \right)}^2}}} = \frac{{3 \cdot 1 + 2 \cdot \frac{1}{2}}}{{1 - \frac{1}{2}}} = ...?\]

OpenStudy (aaronandyson):

3-1/1 - 1/2?

OpenStudy (michele_laino):

no, please: \[\frac{{3{{\left( {\cot \left( A \right)} \right)}^2} + 2{{\left( {\sin \left( A \right)} \right)}^2}}}{{{{\left( {\tan \left( A \right)} \right)}^2} - {{\left( {\cos \left( A \right)} \right)}^2}}} = \frac{{3 \cdot 1 + 2 \cdot \frac{1}{2}}}{{1 - \frac{1}{2}}} = \frac{{3 + 1}}{{\frac{1}{2}}} = \frac{4}{{\frac{1}{2}}} = 4 \cdot 2 = ...?\]

OpenStudy (aaronandyson):

8.

OpenStudy (michele_laino):

that's right!

OpenStudy (aaronandyson):

can you help me with a few more problems?

OpenStudy (michele_laino):

yes! Please wait I have to go to lunch

OpenStudy (aaronandyson):

I'll post a new question and tag you okay??